OpenStudy (anonymous):

Solve for x. |x – 2| + 10 = 12

OpenStudy (anonymous):

just write |x - 2| like x-2 o -x+2

OpenStudy (kirbykirby):

|x-2|=2 so -(x-2)=2 or (x-2)=2